OpenStudy (anonymous):

An initial population of 295 quail increases at an annual rate of 7%. Write an exponential function to model the quail population. What will the approximate population be after 2 years?

OpenStudy (anonymous):

the growth rate is 7% = 0.07 so try \[y = 295(1.07)^{t}\] where t is years and y is the population

OpenStudy (anonymous):

So just simplify 295=(1.07)^2?

OpenStudy (anonymous):

sort of. it's y = 295(1.07)^2

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

generally speaking for these types of problems (growth rates) \[y = a(1 + r)^{t}\] is the formula you use. a is the initial population, and r is the growth rate

OpenStudy (anonymous):

I got 336.7455. Am I correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just round it because the question asks for an approximate. unless its the zombie quail apocalypse where .7455 quails are running around!